A government bond issued in France has a coupon rate of 7 percent, a face value of 100 euros, and matures in five years. The bond pays annual interest payments. Calculate the price of the bond (in euros) if the yield to maturity is 3.5 percent.
Answer #1

Annual coupon=100*7%=7

Hence price of bond=Annual coupon*Present value of annuity factor(3.5%,5)+100*Present value of discounting factor(3.5%,5)

=7*4.515052375+100*0.841973166

=115.80(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=7[1-(1.035)^-5]/0.035

=7*4.515052375

2.Present value of discounting factor=100/1.035^5

=100*0.841973166